MONT 109N – Mathematics Across Cultures Individual Paper 2 – MFA Visit, April 16 due: Wednesday, April 20

Paper Topic

For this paper, aim for about 3 pages, with one or two paragraphs in response to each of the following questions about the objects on display in halls LG32 and LG33 of the Art of the Americas wing.

1. Many of the objects on display are ceramic cups, bowls, and plates. Do we know what purposes those objects served for their owners?

2. The painted illustrations on some of these ceramics record specific events with precise dates. Describe one example of this, the associated date, and the event that is represented.

3. Others show representations of scenes from Maya religious beliefs or the myths that were a large part of their culture (e.g. the epic called the Popol Vuh). Describe one example of that sort.

4. What do these objects tell us about more everyday aspects of the lives of the Maya nobles – what kinds of clothing did they wear? How did they style their hair? What kinds of music might they have made or listened to? What kinds of food and drink did they enjoy? If you know modern Mexican food, can you see any continuities? Along related lines, two examples of current English words that can be traced back to Mayan languages are mentioned in these exhibits. What are they?

5. Many modern viewers of these objects find them strikingly beautiful and wonderful examples of artistic expression. Do you feel that way too? How did the Classic Maya view artistic creativity? And how did that relate to their writing, mathematics, and other parts of their culture? What does the Maya word ts’ib denote?

6. Now that you have seen more of the actual products of Classic Maya culture, have the views you formed from our earlier study of their mathematics and astronomy and the ultimate fate of their civilization changed? If so, in what way? If not, can you say why not? When a human culture (or civilization) collapses, what is lost?

Note: If you are not able to come on the trip, images of many of the objects we will see and some of the explanatory notes are available online at the MFA website:

                  www.mfa.org/search/collections

Enter “Maya ceramics.” There are about 10 pages of 16 images of objects that are accessible this way, including quite a few that are not on display in the museum. Unfortunately, some of the very informative text explanations we will see in the museum is not included in this. However, you should be able to find enough material to answer (at least) parts of 1, 3, parts of 4, 5, 6. To substitute for parts that you cannot answer, pick any one object that is not on display and describe that.